# PySCIPOpt-ML: Embedding Trained Machine Learning Models into Mixed-Integer Programs

### v1

@@ Abstract
A standard tool for modelling real-world optimisation problems is mixed-integer programming (MIP). However, for many of these problems there is either incomplete information describing variable relations, or the relations between variables are highly complex. To overcome both these hurdles, machine learning (ML) models are often used and embedded in the MIP as surrogate models to represent these relations. Due to the large amount of available ML frameworks, formulating ML models into MIPs is highly non-trivial. In this paper we propose a tool for the automatic MIP formulation of trained ML models, allowing easy integration of ML constraints into MIPs. In addition, we introduce a library of MIP instances with embedded ML constraints. The project is available at https://github.com/Opt-Mucca/PySCIPOpt-ML.
@@ 1 Introduction and Related Work
Many problems coming from real-world applications can be modelled using Mixed-Integer Programming (MIP). They can be expressed as
@@
$x_{i}\in\mathbb{Z},\ \forall i\in\mathcal{I},$
where each $f_{i}:\mathbb{R}^{n}\rightarrow\mathbb{R},i=0,\dots,m$ is continuous and $\mathcal{I}\subseteq[n]$ denotes the index set of the integer-constrained variables. Due to its ability to represent complex systems and decision-making processes across various industries, MIP has become a standard tool to solve optimisation problems arising in real-world applications.
A common challenge of finding a suitable MIP formulation (P) of a system is the presence of unknown or highly complex relationships. This poses a problem since traditional MIP approaches rely on precise definitions of constraints and the objective function. To address this challenge, Machine Learning (ML) predictors are often used to approximate these relationships and are then embedded in (P) as surrogate models. To embed a trained ML predictor $g:\mathbb{R}^{n}\rightarrow\mathbb{R}$, it must first be formulated as a MIP, before it can be added to (P). However, with the growing popularity and accessibility of different frameworks [1, 2, 3, 4, 5], automatically formulating trained ML models coming from these different architectures into MIPs has become a highly non-trivial, yet necessary task.
@@
In this paper, we present PySCIPOpt-ML, an open-source tool for the automatic formulation of trained ML predictors as MIPs. By directly interfacing with a broad range of commonly used ML frameworks, PySCIPOpt-ML allows for the easy integration of ML constraints into MIPs, which can then be solved by the state-of-the-art open-source solver SCIP [6]. The package supports various models from Scikit-Learn [2], XGBoost [3], LightGBM [4], and PyTorch [5], making it a general tool to easily optimise MIPs with embedded ML constraints (see Figure 1). To represent predictors as MIPs, we use formulations that prioritise numerical stability to enhance solution robustness. In addition, we introduce an expandable library of MIP instances with embedded ML constraints based on real-world data.
@@
    We utilise MIP formulations prioritising numerical stability, especially when solving with SCIP (see Section 3).
@@
In recent years, more and more industries have leveraged ML frameworks to approximate unknown or complex relationships in optimisation problems. For instance, ML predictors have been used to optimise cancer treatment plans [7], minimise energy consumption [8], and in the design of energy systems [9]. Embedded ML predictors have also been used in the optimisation of gas production and water management systems [10], as well as in the optimisation of retail prices [11] and locations [12], to name only a few examples. For surveys of embedded ML models through surrogate optimisation, see [13, 14, 15].
The increasing demand for embedding ML predictors into MIPs has motivated the development of various tools for automatic integration. The commercial MIP solver Gurobi released the Python package Gurobi Machine Learning [^3] https://github.com/Gurobi/gurobi-machinelearning [16], which integrates trained regression models into optimisation problems that can then be solved with Gurobi. Another package that relies on Gurobi is Janos [17], which supports linear and logistic regression, as well as neural networks with rectified linear unit (ReLU) activation functions. ENTMOOT [18] focuses on representing gradient-boosted tree models from LightGBM. On the other hand, reluMIP [19] allows the user to integrate ReLU neural networks trained with TensorFlow [1] within optimisation problems modelled with either Pyomo [20] or Gurobi. OMLT [^4] https://github.com/cog-imperial/OMLT [21] focuses on embedding neural networks and gradient-boosted trees via the general MIP modelling framework Pyomo [20] and the general purpose ML interface ONNX [^5] https://github.com/onnx/onnx.
While many of these packages focus on supporting specific types of predictors, PySCIPOpt-ML allows the user to embed a wide variety of predictors. Specifically, for both regression and classification tasks, neural networks with various activation functions, linear and logistic regression models, decision trees, gradient boosted trees, and random forests.
@@ 2 About the Package
PySCIPOpt-ML directly interfaces with the popular ML frameworks Scikit-Learn [2], XGBoost [3], LightGBM [4], and PyTorch [5]. Each framework allows the user to train ML predictors of different types, all of which can be embedded into a MIP by our package. A detailed overview of what PySCIPOpt-ML supports is given in Table 1.
ML Predictor PyTorch Scikit-Learn LightGBM XGBoost
Decision Trees - R+C R+C R+C
Gradient Boosted Trees - R+C R+C R+C
Random Forests - R+C R+C R+C
Neural Networks R+C R+C - -
Linear Regression - R - -
Logistic Regression - R+C - -
Partial Least Squares - R - -
Table 1: Overview of ML predictors coming from different ML frameworks supported by PySCIPOpt-ML. “R” indicates that (multi-)regression is supported and “C” denotes the support of (multi-)classification. “-” show predictors not supported by a framework and thus not supported by PySCIPOpt-ML.
An advantage of PySCIPOpt-ML is that it seamlessly joins all parts of the modelling, training, and optimisation processes. At no point is there a need to export the ML predictor or MIP model into some intermediary language. A central function (add_predictor_constr()) acts as the main interface between the ML frameworks training the predictor and the solver SCIP solving the problem. To create and solve a MIP problem with an embedded ML predictor, the user must only do the following:
@@
    Define the optimisation problem, scip_model, by first adding the explicit constraints and objective function.
@@
    Given data (input_data,output_data), choose which predictor should be trained and embed the trained predictor automatically in the optimisation problem as a constraint. This is done by simply calling:
@@ 3 MIP Formulations of ML Predictors
To embed an ML predictor in a MIP, we first need to represent the predictor itself as a MIP. The formulation of the predictor has a significant effect on the ability to efficiently optimise the larger optimisation problem. There has been a substantial amount of work on determining the best performing formulations for popular ML predictors. In particular, neural networks with ReLU activation functions [22, 23, 24, 25] and ensemble based methods that use decision trees [26, 18, 27, 16], have received a lot of attention in the past. Often, the most natural formulation of these predictors involves quadratic constraints. Due to performance issues, however, formulations that utilise big-M, indicator, and SOS constraints are in general preferred.
@@ 3.1 Neural Networks
The current standard MIP formulation for feed forward neural networks with ReLU activation functions is to use big-M constraints [22, 24, 21]. These formulations are prone to numerical issues when the bounds on the input are large, and thus benefit greatly from tighter bounds. For this reason, either feasibility based or optimality based bound tightening is typically performed prior to optimisation [22, 21]. It is therefore important that the user specifies bounds on the variables of the wider optimisation problem. Since this is not always possible, PySCIPOpt-ML utilises a SOS1-based formulation to represent ReLU activation functions. By doing so, we aim to maintain numerical stability even in the absence of tight variable bounds, as the validity of the constraints can be enforced by branching [28].
Consider a single neuron in a fully connected layer $k$. Let $x\in\mathbb{R}^{n}$ be the output of the previous layer $k-1$, $y\in\mathbb{R}$ be the output of the given neuron, $w\in\mathbb{R}^{n}$ and $b\in\mathbb{R}$ be the weights and bias connecting the previous layer to the neuron, and $s\in\mathbb{R}_{\geq 0}$ be some slack variable. We use the following formulation of the ReLU activation at the neuron:
@@
As PySCIPOpt-ML uses SCIP [6] as a solver, other activation functions such as sigmoid and tanh can can be added directly.
3.2 Tree-Based Models
To model decision trees, PySCIPOpt-ML uses the formulation first introduced in Gurobi Machine Learning [16]. As opposed to quadratic formulations [26] and linear big-M formulations [27], the Gurobi Machine Learning formulation uses a series of indicator constraints for each leaf node ensuring that any input maps to the appropriate leaf node of the tree. A more detailed overview of this formulation is available at [^6] https://gurobi-machinelearning.readthedocs.io/en/stable/user/mip-models.html. The MIP formulations of gradient boosted trees and random forests are simply a linear combination of the individual decision tree formulations.
@@ 3.3 Argmax Formulation
For many classification tasks, the ML predictor’s evaluation is determined by an argmax function call. That is, if each class $j$ from the index set $\mathcal{J}$ has some predicted score $y_{j}\in\mathbb{R}$, the predictor outputs the class with the highest score. To model this functionality with MIP, we define a variable $m\in\mathbb{R}$, binary variables $z_{j}\in\{0,1\},j\in\mathcal{J}$, and slack variables $s_{j}\geq 0,j\in\mathcal{J}$. The formulation is:
@@
$\sum_{i=1}^{|\mathcal{J}|}z_{j}=1$
In the above formulation, the variable $m$ implicitly represents max$\{y_{j}:j\in\mathcal{J}\}$: If $m$ is less than the maximum, there exists a $j^{\prime}\in\mathcal{J}$ s.t. $y_{j^{\prime}}\geq m$, and for which $y_{j^{\prime}}+s_{j^{\prime}}-m=0$ cannot be valid. On the other hand, if $m$ is greater than the maximum, then all slack variables are non-zero, and $\sum_{i=1}^{|\mathcal{J}|}z_{j}=1$ cannot be satisfied due to the SOS1 constraints. Because $m$ is the maximum, exactly one $z_{j^{\prime}}$ is set to 1, where $y_{j^{\prime}}=\text{max}\{y_{j}:j\in\mathcal{J}\}$.
To truly return the argmax instead of a binary variable corresponding to the location of the maximum value, one can introduce a variable $a\in\mathbb{Z}$, and the constraint:
$a=\sum_{i=1}^{|\mathcal{J}|}j*z_{j}.$
The variable $a$ will then have value corresponding to that returned by the traditional argmax function.

### v2

@@ Abstract
A standard tool for modelling real-world optimisation problems is mixed-integer programming (MIP). However, for many of these problems, information about the relationships between variables is either incomplete or highly complex, making it difficult or even impossible to model the problem directly. To overcome these hurdles, machine learning (ML) predictors are often used to represent these relationships and are then embedded in the MIP as surrogate models. Due to the large amount of available ML frameworks and the complexity of many ML predictors, formulating such predictors into MIPs is a highly non-trivial task. In this paper, we introduce PySCIPOpt-ML, an open-source tool for the automatic formulation and embedding of trained ML predictors into MIPs. By directly interfacing with a broad range of commonly used ML frameworks and an open-source MIP solver, PySCIPOpt-ML provides a way to easily integrate ML constraints into optimisation problems. Alongside PySCIPOpt-ML, we introduce, SurrogateLIB, a library of MIP instances with embedded ML constraints, and present computational results over SurrogateLIB, providing intuition on the scale of ML predictors that can be practically embedded. The project is available at https://github.com/Opt-Mucca/PySCIPOpt-ML.
@@ 1 Introduction and Related Work
Many problems coming from real-world applications can be modelled using Mixed-Integer Programming (MIP), which can be expressed as follows:
@@
$x_{i}\in\mathbb{Z},\ \forall i\in\mathcal{I}.$
Each function $f_{i}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ for $i\in[m]$ is continuous, and $\mathcal{I}\subseteq[n]$ denotes the index set of the integer-constrained variables. Due to a MIP’s ability to represent complex systems and decision-making processes across various industries, MIP has become a standard tool to solve optimisation problems arising in real-world applications [1, 2].
A common challenge of finding a suitable MIP formulation (P) of a system is the presence of unknown or highly complex relationships. This poses a problem since traditional MIP approaches rely on precise definitions of constraints and the objective function. To address this challenge, Machine Learning (ML) predictors are often used to approximate these relationships and are then embedded in (P) as surrogate models. To embed a trained ML predictor $g:\mathbb{R}^{n}\rightarrow\mathbb{R}$, it must first be formulated as a series of MIP constraints before it can be added to (P). However, with the growing popularity and accessibility of different ML frameworks [3, 4, 5, 6, 7], automatically formulating trained ML predictors coming from these different architectures into MIPs has become a highly non-trivial, yet necessary task.
@@
In this paper, we introduce PySCIPOpt-ML, an open-source tool for the automatic formulation and embedding of trained ML predictors into MIPs. By directly interfacing with a broad range of commonly used ML frameworks, PySCIPOpt-ML allows for the easy integration of ML constraints into MIPs, which can then be solved by the state-of-the-art open-source solver SCIP [8]. The package supports various models from Scikit-Learn [4], XGBoost [5], LightGBM [6], Keras [9], and PyTorch [7], making it a general tool to easily optimise MIPs with embedded ML constraints (see Figure 1). To represent predictors as MIPs, we use formulations based on SOS1 constraints [10], which admit valid formulations in the absence of user specified bounds. Alongside PySCIPOpt-ML, we introduce SurrogateLIB, an expandable library of MIP instances with embedded ML constraints based on real-world data. Computational results over this library of instances are presented, aiming to provide intuition on the scale of ML predictors that can be practically embedded.
@@
    We utilise MIP formulations prioritising practicality, since they do not rely on user defined bounds or expensive bound propagation techniques (see Section 3)
@@
  4. 4.
    We present computational experiments showcasing the scale of predictors that can be practically embedded (see Section 5).
@@
In recent years, more and more industries have leveraged ML frameworks to approximate unknown or complex relationships in optimisation problems. For instance, ML predictors have been used to optimise cancer treatment plans [11], minimise energy consumption [12], and in the design of energy systems [13]. Embedded ML predictors have also been used in the optimisation of gas production and water management systems [14], as well as in the optimisation of retail prices [15] and locations [16], to name only a few examples. For surveys of embedded ML predictors through surrogate optimisation, see [17, 18, 19].
The increasing demand for embedding ML predictors into MIPs has motivated the development of various tools for their automatic integration. The commercial MIP solver Gurobi [20] released the Python package Gurobi Machine Learning [21], which integrates trained regression models into optimisation problems that can then be solved with Gurobi. Other packages that rely on Gurobi are OptiCL [22], which supports many Scikit-Learn predictors [4], and Janos [23], which supports linear and logistic regression, as well as neural networks with rectified linear unit (ReLU) activation functions. EML [24] in contrast uses CPLEX [25], and enables the embedding of decision trees and neural networks through the Keras API [9]. ENTMOOT [26] focuses on representing gradient-boosted tree models from LightGBM [6]. On the other hand, reluMIP [27] allows the user to integrate ReLU neural networks trained with TensorFlow [3] within optimisation problems modelled with either Pyomo [28] or Gurobi. Finally, OMLT [29, 30, 31] focuses on embedding a wide variety of neural networks and gradient-boosted trees via the general MIP modelling framework Pyomo [28] and the general purpose ML interface ONNX [^3] https://github.com/onnx/onnx.
While many of these packages focus on supporting specific types of predictors, PySCIPOpt-ML allows the user to embed a wide variety of predictors. Specifically, for both regression and classification tasks: neural networks with various activation functions, linear and logistic regression models, decision trees, gradient boosted trees, random forests, support vector machines, and centroid clustering. In addition, unlike the other packages, PySCIPOpt-ML interfaces directly with an open-source MIP solver. This allows users to dynamically and transparently improve the solving process, e.g., by dynamic bound tightening [32].
@@ 2 About the Package
PySCIPOpt-ML directly interfaces with the popular ML frameworks Scikit-Learn [4], XGBoost [5], LightGBM [6], Keras [9], and PyTorch [7]. Each framework allows the user to train ML predictors of different types, all of which can then be embedded into a MIP by our package. A detailed overview of what PySCIPOpt-ML supports is given in Table 1.
ML Predictor PyTorch Tensorflow / Keras Scikit-Learn LightGBM XGBoost
Decision Trees - -
SCIP/EML/GRB/
OCL/OMLT
SCIP/EML/GRB/
OMLT
SCIP/GRB/OMLT
Gradient Boosted Trees - -
SCIP/GRB/OCL/
OMLT
SCIP/ENT/GRB/
OMLT
SCIP/GRB/OMLT
Random Forests - -
SCIP/GRB/OCL/
OMLT
SCIP/ENT/GRB/
OMLT
SCIP/GRB/OMLT
Neural Networks (ReLU) SCIP/GRB/OMLT
SCIP/EML/GRB/
OMLT/RMIP
SCIP/GRB/JAN/
OCL/OMLT
Neural Networks (other) SCIP/OMLT SCIP/EML/OMLT SCIP/OMLT - -
Linear Regression - -
SCIP/GRB/JAN/
OCL
Logistic Regression - -
SCIP/GRB/JAN/
OCL
Support Vector Machines - - SCIP/OCL - -
Centroid Clustering - - SCIP - -
Table 1: Overview of ML predictors supported by each automatic embedding framework. The following abbreviations are used: SCIP is PySCIPOpt-ML, EML is Emperical Machine Learning, ENT is ENTMOOT, GRB is Gurobi Machine Learning, JAN is Janos, OCL is OptiCL, OMLT is OMLT, and RMIP is reluMIP. Predictor-framework pairs that do not exist or are not supported by any tool are indicated by "-".
*For all supported frameworks and models PySCIPOpt-ML offers multi-regression and multi-classification.
An advantage of PySCIPOpt-ML is that it seamlessly joins all parts of the modelling, training, and optimisation processes. At no point is there a need to export the ML predictor or MIP model into some intermediary language. A central function (add_predictor_constr) acts as the main interface between the ML frameworks training the predictor and the solver SCIP solving the problem. To create and solve a MIP problem with an embedded ML predictor, the user must only do the following:
@@
    Train the ML predictor, predictor, using one of the supported ML frameworks.
@@
    Define the optimisation problem, scip_model. This involves adding the constraints and objective not directly related to the ML predictor.
  3. 3.
    Create variables that represent (input_data,output_data). The output variables are optional as they can be inferred by the ML predictor. Embed the predictor into the MIP by simply calling:
@@
Figure 2: The workflow when embedding ML predictors into a MIP with PySCIPOpt-ML.
@@ 3 MIP Formulations of ML Predictors
To embed an ML predictor in a MIP, we first need to represent the predictor itself as a MIP. The formulation of the predictor has a significant effect on the ability to efficiently optimise the larger optimisation problem. There has been a substantial amount of work on determining the best performing formulations for popular ML predictors. In particular, neural networks with ReLU activation functions [33, 34, 35, 36] and ensemble based methods that use decision trees [37, 26, 30, 21], have received a lot of attention in the past. Often, the most natural formulation of these predictors involves quadratic constraints. Due to performance issues, however, formulations that utilise big-M, and indicator constraints are in general preferred. In the remainder of this section, we will give an overview of how PySCIPOpt-ML models different ML predictors as MIPs.
@@ 3.1 Neural Networks
The current standard MIP formulation for feed forward neural networks with ReLU activation functions uses big-M constraints [33, 35, 29, 38, 39]. Consider a single neuron in a fully connected hidden-layer $k$, where $k\in\mathbb{N}$. Let $x\in\mathbb{R}^{n}$ be the output of the previous layer $k-1$, $y\in\mathbb{R}_{\geq 0}$ be the output of the given neuron, and $w\in\mathbb{R}^{n}$ and $b\in\mathbb{R}$ be the weights and bias connecting the previous layer to the neuron. The standard big-M formulation requires bounds $l,u\in\mathbb{R}$ on the neuron’s output, and a binary variable for the neuron, namely $z\in\{0,1\}$. The formulation is defined as:
$y$ $\geq 0,$
$y$ $\geq\sum_{i=1}^{n}w_{i}x_{i}+b,$
$y$ $\leq\sum_{i=1}^{n}w_{i}x_{i}+b-l(1-z),$
$y$ $\geq uz.$
Big-M formulations are prone to numerical issues and increased running time when the bounds on the input are large, and thus benefit greatly from tighter bounds [40, 41]. For this reason, when optimising MIPs with embedded neural networks, either feasibility based or optimality based bound tightening is typically performed prior to optimisation [33, 29]. It is therefore important that the user specifies bounds on the variables of the wider optimisation problem. Since this is not always possible, PySCIPOpt-ML gives the user the option to decide between the classical big-M formulation or a novel SOS1-based formulation to represent ReLU activation functions. By doing the latter, we aim to maintain a valid formulation even in the absence of variable bounds, and minimise the pegging of solver performance to bound tightening procedures in the case of provided bounds. The downside of the SOS1-based formulation is that the SOS-constraint does not feature in the LP relaxation, and must be enforced by branching [10].
Our SOS1 formulation for an intermediate neuron with ReLU activation function introduces $s\in\mathbb{R}_{\geq 0}$ as a slack variable for the neuron. The formulation is defined as:
@@
To illustrate the advantages of the SOS1 formulation, we present an example comparing it to the standard big-M formulation. Assume we have $l=-10^{9}$, $u=10^{9}$, and that the expression $\sum_{i=1}^{n}w_{i}x_{i}+b$ evaluates to 0. We can consider the case that $z=10^{-6}$, which is within its integer feasibility tolerance of 0, and would be valid from the solver’s perspective. Consider the constraints however:
$y$ $\geq 0$
$y$ $\geq 0$
$y$ $\leq 0+10^{9}(1-10^{-6})$
$y$ $\leq 10^{9}*10^{-6}$
In the big-M formulation for this poorly scaled example, the output $y$ is in the range $[-\epsilon,10^{3}+\epsilon]$, where $\epsilon\in[0,10^{-6}]$ is the constraint feasibility tolerance of the MIP solver. Thus, the output of the neural network for the MIP can vary greatly compared to its actual prediction from the same input. Meanwhile, due to our SOS1 formulation containing no big-M constraints, the error of the individual node’s prediction is bounded by the feasibility tolerance value. We note however, that in practice such an occurrence is uncommon, since it occurs only when the LP solution happens to be integer feasible and the binary variable is free. We observed that when solving MIPs with embedded ML predictors using SCIP [8], most binary variables are resolved through branching, with the MIP solver fixing the binary variable and not allowing any epsilon changes to its value.
While we have focused on the ReLU activation function formulation, we note that PySCIPOpt-ML also supports other activation functions such as sigmoid, and tanh. As PySCIPOpt-ML uses SCIP [8] as a solver, these formulations can be added directly as non-linear expressions, requiring no reformulation or linearisation.
3.2 Tree-Based Predictors
To model tree-based predictors like decision trees, gradient boosted trees, or random forests, PySCIPOpt-ML uses the formulation first introduced in Gurobi Machine Learning [21]. As opposed to quadratic formulations [37] and linear big-M formulations [30], the Gurobi Machine Learning formulation uses a series of indicator constraints for each leaf node ensuring that any input maps to the appropriate leaf node of the tree. In what follows, we will give an overview of how decision trees are modelled in PySCIPOpt-ML. Note that the MIP formulations of gradient boosted trees and random forests are simply a linear combination of the individual decision tree formulations.
For a decision tree, we denote the set of leaves as $\mathcal{L}\subseteq[m]$ and the input to the tree as $x\in\mathbb{R}^{n}$, where $m\in\mathbb{N}$ is the number of nodes of the tree and $n\in\mathbb{N}$ is the number of features. Each leaf $l\in\mathcal{L}$ is defined by a number of constraints on the input features of the tree that correspond to the branches taken in the path leading to $l$. We formulate decision trees by introducing one binary decision variable $\delta_{l}\in\{0,1\}$ for each leaf of the tree, which indicates whether the given input maps to $l$ (i.e., $\delta_{l}=1$) or not (i.e., $\delta_{l}=0$). Since in a decision tree exactly one leaf is chosen, we have to add the following constraint:
$\sum_{l\in\mathcal{L}}\delta_{l}=1.$
To ensure that the input vector maps to the correct leaf, however, we need to introduce additional notation and constraints. For a node $v\in[m]\setminus\mathcal{L}$ in the tree, we denote by $i_{v}\in[n]$ the feature used for splitting, i.e., making the decision, and by $\theta_{v}\in\mathbb{R}$ the value at which the split is made. At a leaf $l\in\mathcal{L}$, we have a set $\mathcal{L}_{l}$ of inequalities of the form $x_{i_{v}}\leq\theta_{v}$ corresponding to the left branches leading to $l$ and a set $\mathcal{R}_{l}$ of inequalities of the form $x_{i_{v}}>\theta_{v}$ corresponding to the right branches. For each leaf, the inequalities describing $\mathcal{L}_{l}$ and $\mathcal{R}_{l}$ are imposed with indicator constraints modelling the following relationships:
$\delta_{l}=1\implies$ $x_{i_{v}}\leq\theta_{v}-\frac{\epsilon}{2},\quad\forall x_{i_{v}}\leq\theta_{v}\in\mathcal{L}_{l},$
$\delta_{l}=1\implies$ $x_{i_{v}}\geq\theta_{v}+\frac{\epsilon}{2},\quad\forall x_{i_{v}}>\theta_{v}\in\mathcal{R}_{l}.$
In our implementation, $\epsilon\geq 0$ can be specified by a keyword parameter epsilon in the central function add_predictor_constr for embedding a trained predictor as described in Section 2. By default, PySCIPOpt-ML sets $\epsilon=0$. Note that when $\epsilon$ is smaller than the default feasibility tolerance in SCIP [^4] In SCIP 9.0, the feasibility tolerance is set to $10^{-6}$. (as it is by default), and you have a solution where $|x_{i_{v}}-\theta_{v}|\approx\epsilon$, then SCIP can select an arbitrary child node of that decision in the tree. Note also that larger values for $\epsilon$ risk removing solutions containing values $x_{i_{v}}\approx\theta_{v}$. These removed solutions can change the optimal solution of the instance, or make the instance infeasible.
@@ 3.3 Argmax Formulation
For many classification tasks, the ML predictor’s evaluation is determined by an argmax function call. That is, if each class $j$ from the index set $\mathcal{J}$ has some predicted score $y_{j}\in\mathbb{R}$, the predictor outputs the class with the highest score. To model this functionality with MIP, we define a variable $m\in\mathbb{R}$, binary variables $z_{j}\in\{0,1\},j\in\mathcal{J}$, and slack variables $s_{j}\geq 0,j\in\mathcal{J}$. The formulation we use is then given by:
@@
$\sum_{j=1}^{|\mathcal{J}|}z_{j}=1.$
In the above formulation, the variable $m$ implicitly represents max$\{y_{j}:j\in\mathcal{J}\}$: If $m$ is less than the maximum, there exists a $j^{\prime}\in\mathcal{J}$ such that $y_{j^{\prime}}\geq m$, and for which $y_{j^{\prime}}+s_{j^{\prime}}-m=0$ cannot be valid. On the other hand, if $m$ is greater than the maximum, then all slack variables are non-zero, and $\sum_{i=1}^{|\mathcal{J}|}z_{j}=1$ cannot be satisfied due to the SOS1 constraints. Because $m$ is the maximum, exactly one $z_{j^{\prime}}$ is set to 1, where $y_{j^{\prime}}=\text{max}\{y_{j}:j\in\mathcal{J}\}$.
To truly return the argmax instead of a binary variable corresponding to the location of the maximum value, one can introduce a variable $a\in\mathbb{Z}$, and the corresponding constraint
$a=\sum_{j=1}^{|\mathcal{J}|}j*z_{j}.$